The table below shows the salary for two vice presidents of the United States and also the value of the CPl during a year in which each was vice president. The base year of the CPl is an average of 1982 to 1984 dollars, Place your final answer to the question in SNoodle below and show your work on your scratch paper. CPI (Base Year Year Vice PresidentSalary 35,000 175,400 1982-1984) 0.26 1951 Alen Barkky 2001 What is the nominal percentage change in the vice presidents salary from 1951-2001? What is the real percentage increase in salary of our countrys vice president in 1982-1984 dollars, from 1951-2001?
Answer #1

Nominal Percentage change in vice president's salary

= End value/Beginning value -1

=(175400/35000)-1

=401.14 percentage

Real percentage increase in salary

1951 real income in 1982-1984 base year = 35000/0.26=134615.39

2001 real income in 1982-1984 base year = 175000/01.771=98814.22

So increase in salary = (98814.22/134615.39) -1

= -26.6 percentage